OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -24 and 1536, respectively.

OpenStudy (anonymous):

give me a second for a calculator

OpenStudy (anonymous):

k \[a,ar,ar^2,ar^3,...\] \[a,ar=-24, ar^4=1536\] \[r^3=\frac{1536}{-24}=-64\] \[r=\sqrt[3]{-64}=-4\]

OpenStudy (anonymous):

\[a=\frac{-24}{-4}=6\] so the terms look like \[6\times (-4)^n\]

OpenStudy (anonymous):

an = 6 • (-4)n + 1 an = 6 • 4n an = 6 • 4n - 1 an = 6 • (-4)n - 1

OpenStudy (anonymous):

oh i see they want you to write \[\large a_n=6\times (-4)^{n-1}\]

OpenStudy (anonymous):

Okay I gotcha (:

OpenStudy (anonymous):

lol good!